Exercise 2 Pat and Kris are roommates. They spend most of their time studying (of course), but they leave some time for their favorite activities: making pizza and brewing root beer. Pat takes 4 hours to brew a gallon of root beer and 2 hours to make a pizza. Kris takes 6 hours to brew a gallon of root beer and 4 hours to make a pizza. What is each roommates opportunity cost of making a pizza? Who has the absolute advantage in making pizza? Who has the comparative advantage in making pizza? If Pat and Kris trade foods with each other, who will trade away pizza in exchange for root beer? a. b. c. d. The price of pizza can be expressed in terms of gallons of root beer. What is the highest price at which pizza can be traded that would make both roommates better off? What is the lowest price? Explain
Answer #1

a) Pat can brew 1 gallon of root beer or 2 pizzas in 4 hours which implies that the opportunity cost of brewing 1 gallon of root beer is 2 pizzas. Kris can brew 1 gallon of root beer or 6/4 = 1.5 pizzas in 6 hours which implies that the opportunity cost of brewing 1 gallon of root beer is 1.5 pizzas.

b) Take 12 hours. Pat can brew 12/4 = 3 gallons of root beer or 6 pizzas while Kris can brew 2 gallons of root beer or 3 pizzas. It then appears that Pat has absolute advantage (more production in same labor hours) in brewing root bears as well as in the production of pizzas

c) Trade will be based on comparative advantage (lower opportunity cost). Hence Kris brews root beer and trades it with pizza which is produced by Pat. So Pat trades pizza in exchange for root beer

d) The trade price lies between the opportunity cost. The same for 1 pizza lies between 3/6 = 0.5 gallons of root beer and 2/3 = 0.66 gallons of root beer. Hence, the lowest price for 1 pizza is 0.5 gallons of root beer and the highest price is 0.66 gallons of root beer.